# Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates

### v1

@@ 3.1 Methodology
There are some noteworthy distinctions between the training and estimation phases of our method. The $N$ points used to initially train the surrogate are acquired anywhere in the domain $\mathcal{X}$, and assuming they stem from a contour locating design, they seek to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we will use next (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points used for the hybridization and final surrogate update must originate from $p(\mathbf{x})$, and they are solely exploitative. By nature, they will be clustered around the predicted contour. We contend that having a mix of data from these two realms is superior to either extreme. To demonstrate superiority over the first extreme, in which the entire budget is relegated to training the surrogate over $\mathcal{X}$ ignorant of $p(\mathbf{x})$, we will offer empirical comparisons in our simulated examples. To justify this claim over the second extreme, in which all training data must stem from $p(\mathbf{x})$, we revisit the Herbie example of Figure 1. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful.
Choosing M.
Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for large $M$, particularly with a heftier surrogate like a Bayesian DGP. Simply put, we would like $M$ to be large enough without being too cumbersome.
To reason through this choice, let’s revisit the unbiased Monte Carlo estimate of Eq. (2). Notice the indicator $\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator. If we knew $\alpha$, we could determine $M$ by thresholding this standard error. For example, to achieve a standard error less than 10% of $\alpha$ requires:
To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million.
In reality, when the true $\alpha$ is unknown, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time. Surrogate predictions are gathered along this incremental process, with $\alpha$ estimated along the way to inform the selection of $M$.
3.2 Synthetic experiments
In this section we provide a variety of synthetic experiments in which the training and estimation budgets, $N$ and $B$ respectively, are pre-determined. We start with three functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 160, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 400, and the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 800. We split the total budget of evaluations evenly between training and estimation. Table 1 summarizes these settings, among others. Functions are observed without noise to mimic a deterministic computer simulation. We define failure thresholds and input distributions for each function, resulting in the failure probabilities reported in Table 1. We choose $M$ based on these $\alpha$ values to keep $\sigma_{\alpha}$ conservatively under 5%. All input distributions are variations/combinations of independent uniform and truncated normal distributions; further details of the functions, thresholds, and distributions are reserved for Supplementary Material.
Function True $\alpha$ Surrogate $d$ $n_{0}$ $N$ $B$ $M$
Herbie $7.533\times 10^{-5}$ GP 2 20 80 80 $3.5\times 10^{7}$
Ishigami $1.904\times 10^{-4}$ GP 3 50 200 200 $1.5\times 10^{7}$
Hartmann $1.001\times 10^{-5}$ GP 6 100 400 400 $1.0\times 10^{8}$
Plateau $8.530\times 10^{-4}$ DGP 4 30 120 120 $3.5\times 10^{7}$
@@ 3.2 Synthetic experiments
Table 1: Simulaton settings. True $\alpha$ values are estimated from 10 billion samples.
We begin with a random Latin hypercube sample (McKay et al.,, 2000) of size $n_{0}$ (chosen to be 25% of the total training budget, see Table 1). We train a GP surrogate using the Scikit-learn package in python (Pedregosa et al.,, 2011). We then conduct a sequential design following the entropy contour location scheme of Cole et al., (2023), until we reach a final design of size $N$. Ultimately, we deploy the trained surrogate, in conjunction with a large sample $X_{M}\sim p(\mathbf{x})$, to estimate $\alpha$ with each of the following methods:
@@
  • hybMC ENT: Our hybrid MC estimate (Eq. 7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy. [For now, ignore *hybMC ENT, we will return to it in Section 4.]
  • GP MC $N$: Surrogate MC estimate (Eq. 4) using the surrogate trained on $N$ points.
  • GP MC $N+B$: Surrogate MC estimate (Eq. 4) using a surrogate trained on $N+B$ points.
@@
  • hybMC CLOSE: Inspired by Li and Xiu, (2010), the hybrid MC estimate (Eq. 7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{N}(\mathbf{x}_{i})-t|$).
@@
  • SIIS: Importance sampling estimate (Eq. 6; Peherstorfer et al.,, 2016) with a GMM bias distribution, also fit with the Scikit-learn package. The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{N}(\mathbf{x}_{i})>t\right\}$.
@@
  • SIIS UCB: Same as SIIS, but the bias distribution is trained on the samples whose upper 90% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{N}(\mathbf{x}_{i})+1.645\sigma_{N}(\mathbf{x}_{i})>t\right\}$.
Note, the first four methods are variations of Monte Carlo, and the last two are variations of importance sampling. Crucially, each method uses the exact same trained surrogate (besides GP MC $N+B$, which is trained further) and the exact same $X_{M}$ sample. Differences in performance are thus attributed entirely to the estimation method. The GP MC $N$ method is the only one which does not utilize the entire budget; it serves as a snapshot of the surrogate’s accuracy before it is combined with a hybrid calculation or importance sampling. We repeat each exercise 30 times with re-randomized starting designs.
To demonstrate functionality with a DGP, we additionally consider the 4d plateau function (Booth et al.,, 2024b). This function is known to be non-stationary; it is characterized by flat regions with a steep sloping drop between them. To address the complexity in the response surface, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b) fit with the deepgp R-package (Booth,, 2023). We train each DGP using the contour location scheme of Booth et al., (2024b), which involves proposing triangulation candidates (Gramacy et al.,, 2022) and selecting the candidate on the Pareto front of entropy and uncertainty. To keep all of the code for this example in R for easy integration with the DGP, we leverage the mclust R-package (Scrucca et al.,, 2016) to train the GMM bias distribution. Given the increased complexity of the surrogate, we perform 20 re-randomized repetitions. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
Results are shown in Figure 3, with true failure probability indicated by the red dashed line. The grey shaded region highlights the interval $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the provided $M$. This region provides important context – it captures the intrinsic variability that would accompany any MC estimator. In other words, if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (Eq. 2), we would obtain a boxplot roughly spanning the grey interval. We thus consider any results within this region to be effective estimates. We could tighten the interval (and the boxplots of the MC methods) by increasing $M$ without requiring any additional simulator evaluations, another advantage MC holds over IS.
@@
Figure 3: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions (20 for Plateau).
Our hybMC ENT method (teal) performed well across the board. Unsurprisingly, it was able to improve upon the accuracy of the original surrogate (purple) by leveraging the additional budget of $B$ evaluations. More importantly, for all but the simplest Herbie function, it outperformed the surrogate MC estimates which made use of the full budget (blue). Recall, the (D)GP MC methods acquire points from $\mathcal{X}$, while the hybMC method acquires points from $p(\mathbf{x})$ (after the original training phase). The fact that hybMC is able to outperform the fully-trained surrogate indicates that at some point, it is better to prioritize expensive simulator evaluations for uncertain samples from $p(\mathbf{x})$ than to continue training the surrogate.
The strategy of Li and Xiu, (2010) in allotting simulator evaluations based on predicted distance to the contour (hybMC CLOSE, orange) did not significantly improve upon the accuracy of the surrogate alone. This idea fails to leverage the UQ of the surrogate. It is possible for the surrogate to have higher uncertainty in its pass/fail prediction for a point further from the contour. For example, refer back to Figure 2. The entropy surrounding $x_{1}=1$ is far greater than that surrounding $x_{1}=0.5$, even though the latter of these is predicted to be closer to the contour (blue dashed line). Reserving true simulator evaluation for the $x_{1}=1$ location, where the surrogate would misclassify a failure, is preferable.
The SIIS methods (green and pink) had the poorest performance across the board. They particularly struggled in the simpler settings (Herbie and Ishigami), underperforming the simple GP MC $N$ method which only used half the amount of simulator evaluations. We suspect the small budgets are to blame here; IS just needs more data, particularly when the GMM bias density is far from the ideal. Consequently, these IS variations had their best performance with the Hartmann function where 400 samples were budgeted for estimation. There appears to be very little benefit to the more conservative SIIS UCB (pink) over the traditional SIIS (green).
4 Allocating the simulation budget
Let $T=N+B$ represent the total budget of simulator evaluations, which is typically determined by the computational cost of the simulator. For expensive simulators which require long compute times, $T$ may be only several hundred. We presume $T$ is fixed, and we now address how to split it between $N$ and $B$. In our previous studies, we evenly split $T$ between training the surrogate and estimating $\alpha$, i.e., $N=B=\frac{1}{2}T$. But this decision was admittedly arbitrary. In this section, we propose a scheme for allocating $T$ between $N$ and $B$, specifically for use within our hybrid Monte Carlo estimation scheme. Since the sum of $N$ and $B$ is fixed, this choice boils down to selecting $N$; $B=T-N$ then follows directly. Ideally, we would like to halt training of the surrogate once learning has plateaued, saving the rest of the budget for the hybrid Monte Carlo estimate. If we stop training too early, the surrogate might be too inaccurate for the hybrid MC estimate to overcome. If we continue training for too long, we will not have any samples left for the hybrid estimation procedure. As we saw in Figure 3, in all but the simplest case, a MC estimate without any hybridization is suboptimal.
Figure 4: Left: Surrogate MC failure probabilty estimate for the Ishigami function, updated every 10th acquisition with true $\alpha$ in dashed red. Center/Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. The blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
To monitor the surrogate’s learning, we propose tracking the progress of the surrogate MC estimate, $\hat{\alpha}_{\textrm{SURR}}$ (Eq. 4), which we will now denote as simply $\hat{\alpha}$. This estimator relies only on the trained surrogate and does not require additional simulator evaluations. As a demonstration, the left panel of Figure 4 reports the progress in $\hat{\alpha}$ for the 3d Ishigami function following the setup described in Section 3.2. To avoid excessive calculations, we evaluate the estimate after every 10th acquisision. It is apparent that the surrogate’s learning of the failure contour saturates well before it reaches $N=200$, which we used in our previous experiments. The center panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero (this is more apparent in the right panel which is zoomed in).
How small of an update then is small enough to indicate learning has plateaued? The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$ and $M$. Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$. To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update. This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate. The green dashed lines in the center and right panels of Figure 5 show ±$\pm\hat{\sigma}_{\alpha}$ for the estimated failure probabilities. At the start, updates to $\hat{\alpha}$ are far beyond one standard error. But after a while, the updates sit comfortably within these bounds. We posit that, for a fixed sample $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we halt surrogate training once we have observed two successive updates within ±$\pm\hat{\sigma}$.
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as is the case in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted. If the contour location design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for higher dimensional problems. In our synthetic experiments, we used a minimum of 10 failures for the Herbie, Ishigami, and plateau functions but bumped this up to 40 for the larger dimension of the Hartmann function. Once the surrogate has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. Once this condition has been satisfied twice in a row, halt the sequential design and proceed to the hybrid Monte Carlo estimation following Section 3.1.
This stopping criterion requires the same $X_{M}$ sample be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there will naturally be more variability among the $\hat{\alpha}$’s, and the ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased. Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
Figure 5: Distribution of the chosen $N$ for each repetition of the synthetic examples from Figure 3. Red star indicates the value that was used previously.
Revisiting synthetic experiments.
We implemented this stopping criteria with each of the four functions described in Section 3.2. We used the same total budgets ($T=\{160,400,800,240\}$, respectively), but chose $N$ based on this decision rule. Once $N$ was determined, we set $B=T-N$. All other simulation settings were the same. Figure 5 shows the distribution of chosen $N$ values. The red star indicates the value used previously. Most repetitions halted surrogate training before the preset value, allotting more runs for estimation, but some proceeded longer. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design. The resulting failure probability estimates are shown as the *hybMC ENT method (yellow) in Figure 3. The original hybMC ENT methods already performed well, and we were able to match this performance with the budgeted *hybMC ENT method. We acknowledge the benefit of the budgeting method is not in beating the un-budgeted performance on synthetic problems, but in informing budget allocations in real-world scenarios when there is no clear choice for $N$ and $B$.

### v2

@@ 3.3 Summary and related work
There are some noteworthy distinctions between these two design stages. To illustrate, refer back to the black circles ($X_{n}$) and blue triangles ($X_{M}^{B}$) in the right panel of Figure 3. In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative. By nature, they will be clustered around the predicted contour. We contend that having a mix of data from these two realms is superior to either extreme.
Some of the works we referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings. The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user specified threshold. While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided. If $\epsilon$ is too high, contour location will be halted too soon; if too low, contour location will continue past the point of utility. Our proposed stopping criterion avoids this choice and may be easily automated without problem-specific tuning. In our synthetic experiments, we find that it consistently results in appropriate termination with budget to spare.
Hybrid MC estimators (similar to Eq. 7) have been entertained in other settings. Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples delegated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$. This work, however, does not consider surrogate design and does not leverage surrogate UQ. Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$. Echard et al., (2011) similarly train only on a subset of $X_{M}$. We attempted to recreate the approach of Fuhrländer and Schöps, (2020) in our setting: (i) sample large $X_{M}$, (ii) train an initial surrogate on a random selection of these samples, (iii) use the surrogate with a metric such as entropy to select additional batches from $X_{M}$ for evaluation through the simulator, (iv) update the surrogate, and (v) repeat steps (iii)-(iv) until the budget is exhausted. Yet we were unable to obtain even reasonable results because initial surrogate fits were very poor, and surrogate-informed acquisitions were unable to recover. Random selections from $X_{M}$ were overwhelmingly in high probability regions, not low probability failure regions. Hristov et al., (2019) noticed this conundrum and stated it well: “using a surrogate that is built on frequent events to predict and analyze rare events” can be futile.
To visualize this conundrum in a simple setting, we revisit the Herbie example of Figure 1. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful. Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024b).
4 Synthetic experiments
In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments. We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024b) with a total budget of 200. The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023). The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them. To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2023), with the contour location scheme of Booth et al., (2024b). With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
Function True $\alpha$ $d$ $n_{0}$ $N$ $M$ Surrogate CL Scheme
Herbie $7.533\times 10^{-5}$ 2 20 150 $3.5\times 10^{7}$ GP Cole et al., (2023)
Ishigami $1.904\times 10^{-4}$ 3 50 300 $1.5\times 10^{7}$ GP Cole et al., (2023)
Hartmann $1.001\times 10^{-5}$ 6 100 600 $1.0\times 10^{8}$ GP Cole et al., (2023)
Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024b)
@@ 4 Synthetic experiments
Table 1: Simulation settings. True $\alpha$ values are estimated from 10 billion samples.
Table 1 summarizes the simulation settings. Functions are observed without noise to mimic a deterministic computer simulation. We define failure thresholds and input distributions for each function, resulting in the failure probabilities reported in Table 1. All input distributions are variations/combinations of independent uniform and truncated normal distributions; further details of the functions, thresholds, and distributions are reserved for Supplementary Material. We choose $M$ for each $\alpha$ to keep $\sigma_{\alpha}$ conservatively under 5%.
Figure 5: Distribution of the chosen $n$ across 30 repetitions.
First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios, starting with an initial LHS sample of size $n_{0}$ (Table 1) and repeating for 30 re-randomized Monte Carlo repetitions. To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (atleast as many contour locating acquisitions as original space-filling observations). Figure 5 shows the distribution of chosen $n$ values. Notice, all trials terminated before the allotted $N$, reserving some evaluations for additional estimation, but there was considerable variability across trials. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, randomized across repetitions), then deploy the following estimation methods:
  • All CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
@@
  • Hybrid Entropy: Our two-stage approach using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
@@
  • Hybrid Proximity: Same as “Hybrid Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$). This is inspired by the work of Li and Xiu, (2010).
@@
  • SIIS: Surrogate-informed importance sampling (6) with a GMM bias distribution (Peherstorfer et al.,, 2016). The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$. For our GP surrogates implemented in python, we use the scikit-learn package (Pedregosa et al.,, 2011) to train GMM bias distributions. For our DGP surrogates implemented in R, we use the mclust package (Scrucca et al.,, 2016).
@@
  • SIIS UCB: Same as “SIIS,” but the bias distribution is trained on the samples whose upper 95% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})+1.645\sigma_{n}(\mathbf{x}_{i})>t\right\}$.
Each method uses the same $X_{M}$ and the same surrogate trained up to point $n$. The distinction is in what is done with the remaining $B$ samples. As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training and estimation (Figure 5). Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods. The “Hybrid Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$. It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the predicted mean. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
@@
Figure 6: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions.
Results are shown in Figure 6. Red dashed lines mark the true failure probabilities, and grey shaded regions highlight $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the true $\alpha$ and chosen $M$. This region provides important context – it captures the intrinsic variability that would accompany any MC estimator. In other words, even if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (2), we would obtain a boxplot roughly spanning the grey interval. We thus consider any results within this region to be effective estimates. We could tighten the interval (and the boxplots of the MC methods) by increasing $M$ without requiring any additional simulator evaluations, another advantage MC holds over IS.
Across the board, both importance sampling based methods (“SIIS” and “SIIS UCB”) suffered from high variability in their failure probability estimates. Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance. The more conservative “SIIS UCB” had no apparant advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings. In all scenarios aside from the simple two-dimensional Herbie function, our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“All CL”). Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to reserving some simulator evaluations for the second stage. Most of the time there was agreement between the two criteria for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance. As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.